Vapour pressure of pure water is 40mm. If a non volatile solute is added to it vapour pressure falls by 4mm. Hence molality of solution is? 

Grade:12th pass

1 Answers

Khimraj
5 years ago
Relative lowering in vapour pressure
(Po – P)/P = 4/40 = 0.1 = xs
So xo = 0.9
m = 1000*xs/xo*Mo  =  1000*0.1/0.9*18 = 6.17 m
Hope it cleras....................